OpenStudy (anonymous):

If point C is the midpoint of , what is the coordinate of point B .b c=(1,4) a. a=(5,5)

Directrix (directrix):

 B and A are endpoints of a segment with coordinates (x,y) and (5,5) respectively. The coordinates of C are the average of the coordinates of B and A. So, (x + 5)/2 = 1 and (y + 4)/2 = 4 x + 5 = 2 y + 5 = 8 x = -3 y = 3 The c(0,0)rdinates of point A are (-3, 3).